OpenStudy (anonymous):

If x=6csc(theta) in the equation sqrt(x^2-36) how can you find sin and cos of this? All assuming 0

OpenStudy (jdoe0001):

is there any context to it? \(\bf \sqrt{6csc^2(\theta)-36}\) doesn't have to have a sine or cosine per se

OpenStudy (anonymous):

Not really, it says "Use the given substitution to express the given radical expression as a trig function without radicals" but it's asking for\[\sin(\theta) and \cos(\theta). \] I just dont know how to find out what theta is from what it has given.

OpenStudy (jdoe0001):

so it needs to be expressed in terms of sine AND/OR cosine?

OpenStudy (anonymous):

The question asks for both sin and cos of theta.

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

not sure I follow.... can you post a quick screenshot of the material?

OpenStudy (jdoe0001):

I can see AND/OR happening.... and once you get either, you can expand to the other

OpenStudy (anonymous):

Sorry for the confusion, I just don't really understand the question? I don't know how to simplify from the original equation to find theta and take that to finding sin/cos

OpenStudy (jdoe0001):

well.... I can see what they're asking... the sine and cosine of the angle. which is in the 1st Quadrant however I doin't see any " = " equals signs there to equate it against we could simplify it to be cosine based or sine based.... but not equate it to any angle

OpenStudy (jdoe0001):

so.. there isn't really any EQUATion

OpenStudy (anonymous):

So how can you simplify the original square root if there isnt any equation? That's what is confusing me so much, I don't know how to solve it when there's nothing to set it equal to...

OpenStudy (jdoe0001):

exactly, there's nothing equals to we can surely simplify it some.... but that won't equate anything, or any angle for that matter

OpenStudy (jdoe0001):

I assume there might have been a =0 somewhere, but since to have been left out, dunno

OpenStudy (anonymous):

What is theta when it equal zero?

OpenStudy (jdoe0001):

well, if use use \(\bf 0=\sqrt{[6csc(\theta)]^2-36}\) we'd be assuming that's so, but we really dunno

OpenStudy (jdoe0001):

if we use I meant anyhow so... is a stretch there..... some = was left out it seems

OpenStudy (anonymous):

apparently this is what they wanted? Not sure how that works out. Thanks for your help!

OpenStudy (jdoe0001):

looks incomplete.... or maybe there's an example somewhere in your material on what they were doing